OpenStudy (vortish):

Find the min y-value on the graph of y = f(x) f(x)=x^2+8x-5

OpenStudy (vortish):

the formula for this is y= ax^2+bx+c , a = 0 x= -b/2a were b=-8 correct

jimthompson5910 (jim_thompson5910):

x= -b/(2a) x= -8/(2(1)) x = -4 now plug this into f(x)=x^2+8x-5 to find the min

OpenStudy (vortish):

i come up with (4)^2+8(-4)-5 which is 16-32-5 which come up to -21

jimthompson5910 (jim_thompson5910):

good

OpenStudy (vortish):

thanks Mr Thompson i been trying to figure this one out for a while

jimthompson5910 (jim_thompson5910):

np

OpenStudy (vortish):

It is nice to have people that can help